OpenStudy (ilovebmth1234):

Find the height, in feet, of the ball after 3 seconds in the air? help?a ball is launched from a sling shot. its height, h(x), can be represented by a quadratic function in terms of time,x,in seconds. After 1second, the ball is 121 feet in the air, after 2 seconds, it is 224 feet in the air.

OpenStudy (lastdaywork):

This is similar to previous question. Assume a general quadratic equation & make three equations with three unknown variables.

OpenStudy (ilovebmth1234):

ok

OpenStudy (lastdaywork):

Tell me the three equations; so that I can check them :)

OpenStudy (lastdaywork):

What was your general quadratic equation ??

OpenStudy (ilovebmth1234):

what do you mean? this? h(x)=ax^2+mx+b?

OpenStudy (lastdaywork):

In that case; h(1) = a(1^2) + m(1) + b = 121

OpenStudy (ilovebmth1234):

h(1)=a1^2+m1+b h(2)=a2^2+m2+b h(3)=a3^2+m3+b

OpenStudy (lastdaywork):

As we don't have a value for h(3) ; we can use it to deduce the function h. Read the question again and try to find one more equation.

OpenStudy (ilovebmth1234):

im not finding it?

OpenStudy (ilovebmth1234):

uggghhh i suck at this /.\

OpenStudy (ilovebmth1234):

is the answer 327?

OpenStudy (lastdaywork):

OS was down for me last night..then I went to sleep.. At time t = 0; distance traveled = 0 ; thus h(0) = a(0^2) + m(0) + b = 0 Hence, b = 0 So we are left with two equations and two unknowns. a(1^2) + m(1) = 121 a(2^2) + m(2) = 224 Find a and m Finally calculate h(3)

OpenStudy (lastdaywork):

32 is NOT the correct answer. See http://www.wolframalpha.com/input/?i=x+%2B+y+%3D+121+%3B+4x+%2B+2y+%3D+224+%3B+9x+%2B+3y+%3D+z

OpenStudy (ilovebmth1234):

god i suck at this

OpenStudy (lastdaywork):

First of all; fan me coz I can't reply your PMs.. Also, can you tell me where you are stuck ??

OpenStudy (ilovebmth1234):

im on a different problem but i dont know what to do after you get your 3 equations

OpenStudy (lastdaywork):

Can you tell me the question and your equations..

OpenStudy (ilovebmth1234):

photo necessities produce camera cases. they have found that the cost, C(x) of making x camera cases is a quadratic function in the terms of x. the company also discovered that it costs $23 to produce 2 cases $103 to produce 4 cases and $631 to produce 10 cases. complete the function that represents the cost, C(x) to produce x camera cases

OpenStudy (ilovebmth1234):

C(x)=2a^2+m2+b C(x)=4a^2+m4+b C(x)=10a^2+m10+b

OpenStudy (lastdaywork):

First, you should explicitly write the function - \[C(x) = ax^2+mx+b\] Now, can you see the (typographical) mistake in your equations. Also, replace C(2), C(4), C(10) with their respective values.. Please, re-write the equations.

OpenStudy (ilovebmth1234):

C(2)=2a^2+m2+b C(4)=4a^2+m4+b C(10)=10a^2+m10+b

OpenStudy (ilovebmth1234):

so whats do i do now?

OpenStudy (lastdaywork):

Actually, the equations should be \[C(2) = a(2^2) + m(2) + b\] \[C(4) = a(4^2) + m(4) + b\] \[C(10) = a(10^2) + m(10) + b\]

OpenStudy (lastdaywork):

We already know that C(2) = 23 C(4) = 103 C(10) = 631 So replace these values in the above equations and tell me what you get.

OpenStudy (ilovebmth1234):

23=a(22)+m(2)+b 103=a(42)+m(4)+b 631=a(102)+m(10)+b

OpenStudy (lastdaywork):

Er.. \[a^2 = a*a\]

OpenStudy (lastdaywork):

So, \[2^2 \neq 22\]

OpenStudy (ilovebmth1234):

thats 4

OpenStudy (lastdaywork):

Yea, can you re-write the equations ??